OpenStudy (anonymous):

Evaluate ∫∫ over D (cos(sqrt(x^2+y^2)) dA by changing polar coordinates where the disk is with center of origin and radius 2

OpenStudy (anonymous):

I got up to ∫∫[0 to 2pi] [0 to 2] cos(sqrt(r)) r dr dt

OpenStudy (anonymous):

but I can't solve for it! How do integrate r*cos(sqrt(r))

OpenStudy (anonymous):

\[\int\limits\int\limits_{D}=\int\limits_{0}^{2\pi} d\theta\ \int\limits_{0}^{2}r \cos rdr\]

OpenStudy (anonymous):

omg you are correct.

OpenStudy (anonymous):

spent literally 30 mins trying to see what I did wrong. That was quick thanks a lot man

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

by the way:int( rcos r) it's done by integration by parts

OpenStudy (anonymous):

can't I do substitution

OpenStudy (anonymous):

you can try

OpenStudy (anonymous):

haha I just did and it didn't work, alright thanks again!

OpenStudy (anonymous):

if you have problems look here, :) http://en.wikipedia.org/wiki/Integration_by_parts

OpenStudy (anonymous):

there exacly you example

OpenStudy (anonymous):

your*

OpenStudy (anonymous):

Yeah I got it now, thanks a lot for your help!